# Pseudoentanglement from tensor networks

### v1

@@ Abstract
Introduction. Entanglement is a fundamental resource in quantum information science and an important principle in theoretical physics, from condensed matter to gravity Horodecki and Horodecki (2009); Kitaev and Preskill (2006); Levin and Wen (2006); Swingle (2012); Maldacena and Susskind (2013); Liu and Suh (2014); Kaufman et al. (2016); Abanin et al. (2019). At the same time, the entanglement structure of large many-body states is very challenging to observe. Specifically, there exist $N$-qubit states with very low entanglement that are computationally indistinguishable (i.e., indistiguishable to any observer whose resources scale polynomially in $N$) from states with much higher entanglement, such as Haar-random states. This property is knowns as pseudoentanglement Aaronson et al. (2023); Giurgica-Tiron and Bouland (2023); Bouland et al. (2023).
@@
Pseudoentangled matrix product states.—We first exhibit our construction on the simplest tensor network: the one-dimensional matrix product state (MPS). A MPS, sketched in Fig. 1(a), is a many-body state whose amplitudes are expressed as $\innerproduct{i_{1},\dots i_{N}}{\psi}={\rm Tr}(A^{[1]}_{i_{1}}\cdots A^{[N]}_{i_{N}})$, with each $A^{[j]}_{i_{j}}$ a $\chi\times\chi$ matrix; $\chi$ is a numerical cut-off known as the bond dimension, which upper-bounds the Schmidt rank of the state about any bond on the 1D line and thus imposes an area-law, Fig. 1(b). A standard route to prepare MPSs in the lab is to use ‘staircase’ circuits  [^2] This approach is also known as “holographic” Foss-Feig et al. (2021); Anand et al. (2023), though not to be confused with the other usage of the word in this paper., where one associates $A_{i,\alpha,\beta}^{[j]}=\bra{i,\beta}U^{[j]}\ket{0,\alpha}$, with $U^{[j]}$ a unitary gate acting on a qubit and a $\chi$-state qudit. This is sketched in Fig. 1(c). In this work we take $\chi=2^{\nu}$, so each $U^{[j]}$ is a $(\nu+1)$-qubit gate. The staircase circuit comprises $L$ gates; for the last gate, all $\nu+1$ output qubits become part of the physical system, which in all is made of $N=L+\nu$ qubits. We take our unitary gates $\{U^{[j]}\}_{j=1}^{L}$ to be drawn from the recently introduced ‘PFC ensemble’ Metger et al. (2024) of pseudorandom unitaries (PRUs). These are efficiently implementable transformations that are computationally indistinguishable from Haar-random unitaries (see  [^3] See Supplementary Information., Metger et al. (2024) for more details). We denote the resulting ensemble of $N$-qubit pure states by $\mathcal{E}_{\rm pMPS}$.
@@
We split this task in three parts. First, by definition of the PFC ensemble Metger et al. (2024), the pseudorandom ensemble $\mathcal{E}_{\rm pMPS}$ is computationally indistinguishable from its counterpart $\mathcal{E}_{\rm pMPS^{\prime}}$ where the pseudorandom PFC gates are replaced by genuinely random permutations, phase, and Clifford gates. Next, we prove that $\mathcal{E}_{\rm pMPS^{\prime}}$ is indistinguishable from its counterpart made of genuine Haar-random gates, which we term $\mathcal{E}_{\rm rMPS}$:
@@
We have $\|\rho^{(m)}_{\mathcal{E}_{\rm pMPS^{\prime}}}-\rho^{(m)}_{\mathcal{E}_{\rm rMPS}}\|_{\rm tr}\leq O(Nm/\sqrt{\chi})$.
@@
Figure 2: Construction of pseudoentangled states bsed on arbitrary isometric tensor networks. (a-c) Isometric realization of a PEPS on a $2D$ square lattice and a holographic tensor network on a $\{6,4\}$ tiling of the hyperbolic plane. Arrows denote the direction of unitarity. Gray triangles in (a) are |$\ket{0}$ states, gray circles in (c) are Bell pair states. Thin lines (in (a) only) are qubits, thick lines are $\chi$-state qudits. (b-d) Partition functions obtained from averaging over Haar-random unitaries. Green and yellow sites represent “bulk” spins ($\tau$) and “boundary” spins ($\sigma$), respectively. Solid bonds give overlaps between permutations, dashed bonds give Weingarten functions. Bonds marked by a $\times$ can be dropped to obtain a spanning tree of the interaction graph, used in our proof of computational pseudorandomness.
Pseudoentanglement in general isometric tensor networks. This construction is straightforwardly generalizable to other geometries beyond one dimension, e.g. to PEPS in 2D Verstraete et al. (2006); Cirac et al. (2021), which can also be realized as isometric tensor networks Zaletel and Pollmann (2020) subject to some constraints on their bond dimensions, see Fig. 2(a-b). We generalize the proof for MPS given above to arbitray isometric tensor networks in Note (2); here we summarize the main idea.
@@
Figure 3: (a) Schematic of minimum cut for the entanglement of a boundary interval $A$ in a holographic tensor network. (b) Scaling of the entanglement entropy in holographic pseudoentangled states (here each $\chi$-state qudit in the output state is split into $\nu=\log_{2}(\chi)$ qubits). A Page-like scaling $\sim|A|\log(2)$ extends up to $|A|\sim\log(\chi)=\omega(\log N)$, then the minimum cut $|\gamma_{A}|$ starts deviating from the volume $|A|$; in hyperbolic tilings one has have $|\gamma_{A}|\sim\log|A|$ on large length scales.
Holographic tensor networks. An important application of our construction is to tensor network models of holography, where a quantum system is viewed as the “boundary” of a higher-dimensional “bulk” space, and its entanglement structure reflects the geometry of the bulk. These models Pastawski et al. (2015); Swingle (2012); Qi (2013); Hayden et al. (2016); Qi et al. (2017); Jahn and Eisert (2021) have attracted considerable interest as toy models of the AdS/CFT correspondence and its links to quantum error correction. They generally feature tensors arranged on a tiling of $2D$ hyperbolic space, which may or may not have physical legs in the bulk, terminating with “dangling” legs on the boundary which represents the physical system of interest. Holographic tensor networks are very well suited to our construction, since they encode the bulk isometrically into the boundary; i.e., the tensor network may be viewed as a circuit propagating information unitarily outward from the center of the hyperbolic space toward its boundary. (We do not consider recently introduced non-isometric codes here, though that would be an interesting direction for future work Akers et al. (2022); DeWolfe and Higginbotham (2023).)
@@
Discussion. We have constructed new models of pseudoentangled quantum state ensembles by combining pseudorandom unitaries and isometric tensor networks. Compared to prior constructions of pseudoentanglement, this approach provides more flexibility in the choice of entanglement structures that may be computationally hidden. These include pseudo-area-law scaling in any dimension as well as holographic entanglement which obeys a Ryu-Takayanagi formula, Eq. (9). The implications of our construction for holography remain to be explored, as well as possible extensions of our construction to more general and non-isometric holographic codes Akers et al. (2022); DeWolfe and Higginbotham (2023).
@@
Note added. While working on this manuscript we became aware of a forthcoming work by Akers, Bouland, Chen, Kohler, Metger and Vazirani Akers et al. (2024) that introduces a different construction of holographic pseudoentangled states. Additionally, during or after completion of this manuscript, several related works appeared on arXiv: Ref. Engelhardt et al. (2024) by Engelhardt et al introduced a different construction of pseudoentangled holographic states; Ref. Schuster et al. (2024) by Schuster, Huang and Haferkamp introduced a construction of pseudoentangled states based on shallow circuits of PFC PRUs—their 1D construction is similar, but not identical, to our MPS construction, and the techniques used in our proofs are different; finally Ref. Lami et al. (2024) numerically studied random MPS and PEPS states in the regime of $\chi={\rm poly}(N)$, finding results that become consistent with ours when extrapolated to $\chi=\omega({\rm poly}(N))$.
Acknowledgments. MI gratefully acknowledges discussions with Adam Bouland, Tudor Giurgica-Tiron, and Nick Hunter-Jones.
@@
S2 Computational indistinguishability between $\mathcal{E}_{\rm pMPS^{\prime}}$ and $\mathcal{E}_{\rm rMPS}$
Here we give a proof of Lemma 1, stating the indistinguishability between the ensemble of MPS made out of Haar-random gates, $\mathcal{E}_{\rm rMPS}$, and the analogous ensemble where the MPS is made out of random ‘PFC’ gates, $\mathcal{E}_{\rm pMPS^{\prime}}$. To this end, we use a key property of the PFC ensemble:
@@
$\|\Phi^{(m)}_{\rm Haar}\otimes\mathcal{I}(\rho)-\Phi^{(m)}_{\rm PFC}\otimes\mathcal{I}(\rho)\|_{\rm tr}\leq O(m/\sqrt{d}),$ (S5)
where $\Phi^{(m)}_{\rm Haar}\otimes\mathcal{I}(\rho)=\mathbb{E}_{U\sim{\rm Haar}}[(U^{\otimes m}\otimes I)\rho(U^{\otimes m}\otimes I)^{\dagger}]$ is the $m$-fold twirling channel for the Haar ensemble (acting only on the $\mathcal{H}_{A}^{\otimes m}$ factor of the Hilbert space), and $\Phi^{(m)}_{\rm PFC}$ is the same for the PFC ensemble.
@@
The moment operators for the two ensembles $\mathcal{E}_{\rm pMPS^{\prime}}$, $\mathcal{E}_{\rm rMPS}$ can be written as
@@
$\rho^{(m)}_{\mathcal{E}_{\rm pMPS^{\prime}}}$ $=\Phi^{(m)}_{\rm PFC}|_{[L,L+\nu]}\circ\Phi^{(m)}_{\rm PFC}|_{[L-1,L-1+\nu]}\circ\cdots\circ\Phi^{(m)}_{\rm PFC}|_{[1,\nu+1]}(\rho_{0}),$ (S7)
@@
$\|\rho^{(m)}_{\mathcal{E}_{\rm rMPS}}-\rho^{(m)}_{\mathcal{E}_{\rm pMPS^{\prime}}}\|_{\rm tr}\leq L\epsilon,$ (S8)
with $\epsilon$ the approximation error of the PFC ensemble on $\nu+1$ qubits: $\epsilon\leq O(m/\sqrt{\chi})$, where $\chi=2^{\nu}$ is the MPS bond dimension. We will prove the statement by induction.
  • ($L=1$): The whole system consists of $\nu+1$ qubits. We have $\|\Phi^{(m)}_{\rm Haar}(\rho_{0})-\Phi^{(m)}_{\rm PFC}(\rho_{0})\|_{\rm tr}\leq\epsilon$.
@@
    $\rho_{2}$ $=\Phi^{(m)}_{\rm PFC}|_{[L-1,L-1+\nu]}\circ\cdots\circ\Phi^{(m)}_{\rm PFC}|_{[1,\nu+1]}(\rho_{0}).$ (S10)
@@
    $\|\Phi^{(m)}_{\rm Haar}|_{[L,L+\nu]}(\rho_{1})-\Phi^{(m)}_{\rm PFC}|_{[L,L+\nu]}(\rho_{2})\|_{\rm tr}$ $\leq\|\Phi^{(m)}_{\rm Haar}|_{[L,L+\nu]}(\rho_{1})-\Phi^{(m)}_{\rm PFC}|_{[L,L+\nu]}(\rho_{1})\|_{\rm tr}$
    $\qquad+\|\Phi^{(m)}_{\rm PFC}|_{[L,L+\nu]}(\rho_{1})-\Phi^{(m)}_{\rm PFC}|_{[L,L+\nu]}(\rho_{2})\|_{\rm tr}.$ (S11)
    The first term on the rhs is $\leq\epsilon$ by direct application of Eq. (S5). The second term, by monotnonicity of the trace distance under quantum channels, is $\|\Phi^{(m)}_{\rm PFC}(\rho_{1}-\rho_{2})\|_{\rm tr}\leq\|\rho_{1}-\rho_{2}\|_{\rm tr}$; but (up to tensoring with a trivial state |$\outerproduct{0}{0}^{\otimes m}$) $\rho_{1}-\rho_{2}$ is the difference between the two moment operators at step $L-1$, which by hypothesis is $\leq(L-1)\epsilon$. The result follows.
We remark that this proof did not use the MPS structure specifically, but only the isometric structure of the tensor network (i.e. the ability to write it as a sequential application of unitary gates). As long as the network consists of ${\rm poly}(N)$ gates, the error incurred by replacing Haar-random gates by PFC gates remains small, $O({\rm poly}(N)m/\sqrt{\chi})$.
@@
Here we modify the proof presented in the main text for MPS to general isometric tensor network architectures.
@@
    Unitary vertices must have both incoming and outgoing edges, and must respect the condition $\prod_{i\in{\rm incoming}}\chi_{i}=\prod_{i\in{\rm outgoing}}\chi_{i}$ (unitarity).
@@
    Input vertices have only outgoing legs. We restrict for simplicity to product states |$\ket{0}$, which have one outgoing edge, and Bell pairs |$\ket{\psi_{\rm Bell}}=(1/\sqrt{\chi})\sum_{j=0}^{\chi}\ket{jj}$, which have two outgoing edges of the same bond dimension $\chi$. Both types of input vertices are illustrated in Fig. 2(a,c).
@@
    Output vertices have only one incoming edge and no outgoing edges. They represent “dangling legs” in standard tensor network notation.
$N$ is the number of output vertices. We assume that the total number of unitary vertices in the graph is $n_{U}\leq O({\rm poly}(N))$. This ensures the states’ efficient preparability.
@@
We can repeat the same argument as in Sec. S2, but instead of $O(N\epsilon)$ as the upper bound we obtain $O(\sum_{u\in{\rm gates}}\epsilon_{u})$, since in this case each gate may act on a Hilbert space of different dimension. We have $\epsilon_{u}\leq O(m^{2}/d_{u})$, with $d_{u}=\prod_{\ell\in{\rm outgoing}(u)}\chi_{\ell}$, and thus
$\|\rho^{(m)}_{\mathcal{E}_{\rm pTNS}}-\rho^{(m)}_{\mathcal{E}_{\rm rTNS}}\|_{\rm tr}\leq O(n_{U}m/\sqrt{d_{\rm min}}),\qquad d_{\rm min}=\min_{u}\prod_{\ell\in{\rm outgoing}(u)}\chi_{\ell}.$ (S12)
Here $n_{U}$ is the total number of unitary gates in the network, again assumed to be $O({\rm poly}(N))$.
@@
We can now perform the Haar averages and obtain a modified graph corresponding to the partition functions in Fig. 2(b,d). The new graph is undirected, weighted and labeled—each edge has one of two labels, that we denote by ‘solid’ and ‘dashed’ in keeping with Fig. 1 and Fig. 2. The new graph is built according to these rules:
  • All edges between two unitary vertices become ‘solid’ and keep their weight.
@@
This graph serves as the interaction graph for the $S_{m}$-valued magnet’s partition function: each vertex hosts a spin $\alpha\in S_{m}$, each edge connecting two spins $\alpha_{i,j}$ contributes a Boltzmann weight $\chi_{\ell}^{|\alpha_{i}\alpha_{j}^{-1}|}$ if it is ‘solid’, $\textsf{Wg}_{d_{u}}(\alpha_{i}\alpha_{j}^{-1})$ if it is ‘dashed’, with $\chi_{\ell}$ and $d_{u}$ the edge weights as defined above. In addition, there is an overall prefactor of $\prod_{\ell\in{\rm Bell}}\chi_{\ell}^{-m}$, with the product running over Bell pair input states in the original graph, coming from the normalization prefactor in |$\outerproduct{\psi_{\rm Bell}}{\psi_{\rm Bell}}^{\otimes m}$.
@@
where we used the results of Sec. S1 to approximate the Weingarten functions, and restricted the product over edges to edges that did not originate from Bell pairs. Now, viewing $d_{u}$ as the product of $\chi_{\ell}$ over outgoing legs of the unitary, we have that each leg connecting two unitaries gives both a factor of $\chi^{+m}$ (from the first term in Eq. (S13)) and a factor of $\chi^{-m}$ (from the second term), which cancel. The only factors to not cancel are $\chi^{-m}$ for legs in the original graph that do not terminate at a unitary. These legs can only terminate at output vertices, so in all the result is $(\prod_{\ell\in{\rm output}}\chi_{\ell})^{-m}=D^{-m}$, with $D$ the physical Hilbert space dimension. The error terms stem only from the Weingarten functions and can be bounded above by $|[1+O(m^{2}/d_{u})]^{n_{U}}-1|$ with $n_{U}$ the number of unitaries. Overall this gives

### v2

@@ Abstract
Introduction.—Entanglement is a fundamental resource in quantum information science and an important principle in theoretical physics, from condensed matter to gravity Horodecki and Horodecki (2009); Kitaev and Preskill (2006); Levin and Wen (2006); Swingle (2012); Maldacena and Susskind (2013); Liu and Suh (2014); Kaufman et al. (2016); Abanin et al. (2019). At the same time, the entanglement structure of large many-body states is very challenging to observe. Specifically, there exist $N$-qubit states with very low entanglement that are computationally indistinguishable (i.e., indistiguishable to any observer whose resources scale polynomially in $N$) from states with much higher entanglement, such as Haar-random states. This property is knowns as pseudoentanglement Aaronson et al. (2023); Giurgica-Tiron and Bouland (2023); Bouland et al. (2023).
@@
Pseudoentangled matrix product states.—We first exhibit our construction on the simplest tensor network: the one-dimensional matrix product state (MPS). A MPS, sketched in Fig. 1(a), is a many-body state whose amplitudes are expressed as $\innerproduct{i_{1},\dots i_{N}}{\psi}={\rm Tr}(A^{[1]}_{i_{1}}\cdots A^{[N]}_{i_{N}})$, with each $A^{[j]}_{i_{j}}$ a $\chi\times\chi$ matrix; $\chi$ is a numerical cut-off known as the bond dimension, which upper-bounds the Schmidt rank of the state about any bond on the 1D line and thus imposes an area-law, Fig. 1(b). A standard route to prepare MPSs in the lab is to use ‘staircase’ circuits  [^2] This approach is also known as “holographic” foss-feig_holographic_2021; anand_holographic_2023, though not to be confused with the other usage of the word in this paper., where one associates $A_{i,\alpha,\beta}^{[j]}=\bra{i,\beta}U^{[j]}\ket{0,\alpha}$, with $U^{[j]}$ a unitary gate acting on a qubit and a $\chi$-state qudit. This is sketched in Fig. 1(c). In this work we take $\chi=2^{\nu}$, so each $U^{[j]}$ is a $(\nu+1)$-qubit gate. The staircase circuit comprises $L$ gates; for the last gate, all $\nu+1$ output qubits become part of the physical system, which in all is made of $N=L+\nu$ qubits. We take our unitary gates $\{U^{[j]}\}_{j=1}^{L}$ to be drawn from the recently introduced ‘PFC ensemble’ Metger et al. (2024) of pseudorandom unitaries (PRUs). These are efficiently implementable transformations that are computationally indistinguishable from Haar-random unitaries (see  [^3] See Supplementary Information., Metger et al. (2024) for more details). We denote the resulting ensemble of $N$-qubit pure states by $\mathcal{E}_{\rm pMPS}$.
@@
We split this task in three parts. First, by definition of the PFC ensemble Metger et al. (2024), the pseudorandom ensemble $\mathcal{E}_{\rm pMPS}$ is computationally indistinguishable from its counterpart $\mathcal{E}_{\rm rMPS^{\prime}}$ where the pseudorandom PFC gates are replaced by genuinely random permutations, phase, and Clifford gates. Next, we prove that $\mathcal{E}_{\rm rMPS^{\prime}}$ is indistinguishable from its counterpart made of genuine Haar-random gates, which we term $\mathcal{E}_{\rm rMPS}$:
@@
We have $\|\rho^{(m)}_{\mathcal{E}_{\rm rMPS^{\prime}}}-\rho^{(m)}_{\mathcal{E}_{\rm rMPS}}\|_{\rm tr}\leq O(Nm/\sqrt{\chi})$.
@@
Figure 2: Construction of pseudoentangled states bsed on arbitrary isometric tensor networks. (a-c) Isometric realization of a PEPS on a $2D$ square lattice and a holographic tensor network on a $\{6,4\}$ tiling of the hyperbolic plane. Arrows denote the direction of unitarity. Gray triangles in (a) are |$\ket{0}$ states, gray circles in (c) are Bell pair states. Thin lines (in (a) only) are qubits, thick lines are $\chi$-state qudits. (b-d) Partition functions obtained from averaging over Haar-random unitaries. Green and yellow sites represent “bulk” spins ($\tau$) and “boundary” spins ($\sigma$), respectively. Solid bonds give overlaps between permutations, dashed bonds give Weingarten functions. Bonds marked by a $\times$ can be dropped to obtain a spanning tree of the interaction graph, used in our proof of indistinguishability.
Pseudoentanglement in general isometric tensor networks.—This construction is straightforwardly generalizable to other geometries beyond one dimension, e.g. to PEPS in 2D Verstraete et al. (2006); Cirac et al. (2021), which can also be realized as isometric tensor networks Zaletel and Pollmann (2020) subject to some constraints on their bond dimensions, see Fig. 2(a-b). We generalize the proof for MPS given above to arbitray isometric tensor networks in Note (2); here we summarize the main idea.
@@
Figure 3: (a) Schematic of minimum cut for the entanglement of a boundary interval $A$ in a holographic tensor network. (b) Scaling of the entanglement entropy in holographic pseudoentangled states (here each $\chi$-state qudit in the output state is split into $\nu=\log_{2}(\chi)$ qubits). A Page-like scaling $\sim|A|\log(2)$ extends up to $|A|\sim\log(\chi)=\omega(\log N)$, then the minimum cut $|\gamma_{A}|$ starts deviating from the volume $|A|$; in hyperbolic tilings one has $|\gamma_{A}|\sim\log|A|$ on large length scales.
Holographic tensor networks.—An important application of our construction is to tensor network models of holography, where a quantum system is viewed as the “boundary” of a higher-dimensional “bulk” space, and its entanglement structure reflects the geometry of the bulk. These models Pastawski et al. (2015); Swingle (2012); Qi (2013); Hayden et al. (2016); Qi et al. (2017); Jahn and Eisert (2021) have attracted considerable interest as toy models of the AdS/CFT correspondence and its links to quantum error correction. They generally feature tensors arranged on a tiling of $2D$ hyperbolic space, which may or may not have physical legs in the bulk, terminating with “dangling” legs on the boundary which represents the physical system of interest. Holographic tensor networks are very well suited to our construction, since they encode the bulk isometrically into the boundary; i.e., the tensor network may be viewed as a circuit propagating information unitarily outward from the center of the hyperbolic space toward its boundary. (We do not consider recently introduced non-isometric codes here, though that would be an interesting direction for future work Akers et al. (2022); DeWolfe and Higginbotham (2023).)
@@
Discussion.—We have constructed new models of pseudoentangled quantum state ensembles by combining pseudorandom unitaries and isometric tensor networks. Compared to prior constructions of pseudoentanglement, this approach provides more flexibility in the choice of entanglement structures that may be computationally hidden. These include pseudo-area-law scaling in any dimension as well as holographic entanglement which obeys a Ryu-Takayanagi formula, Eq. (9). The implications of our construction for holography remain to be explored, as well as possible extensions of our construction to more general and non-isometric holographic codes Akers et al. (2022); DeWolfe and Higginbotham (2023).
@@
Note added.—While working on this manuscript we became aware of a forthcoming work by Akers, Bouland, Chen, Kohler, Metger and Vazirani Akers et al. (2024) that introduces a different construction of holographic pseudoentangled states. Additionally, during or after completion of this manuscript, several related works appeared on arXiv: Ref. Engelhardt et al. (2024) by Engelhardt et al introduced a different construction of pseudoentangled holographic states; Ref. Schuster et al. (2024) by Schuster, Haferkamp and Huang introduced a construction of pseudoentangled states based on shallow circuits of PFC PRUs—their 1D construction is similar, but not identical, to our MPS construction, and the techniques used in our proofs are different; finally Ref. Lami et al. (2024) by Lami, De Nardis and Turkeshi analitically and numerically studied random MPS and PEPS states in the regime of $\chi={\rm poly}(N)$, finding results that become consistent with ours when extrapolated to $\chi=\omega({\rm poly}(N))$.
Acknowledgments.—MI gratefully acknowledges discussions with Adam Bouland, Tudor Giurgica-Tiron, and Nick Hunter-Jones.
@@
S2 Computational indistinguishability between $\mathcal{E}_{\rm rMPS^{\prime}}$ and $\mathcal{E}_{\rm rMPS}$
Here we give a proof of Lemma 1, stating the indistinguishability between the ensemble of MPS made out of Haar-random gates, $\mathcal{E}_{\rm rMPS}$, and the analogous ensemble where the MPS is made out of random ‘PFC’ gates, $\mathcal{E}_{\rm rMPS^{\prime}}$. To this end, we use a key property of the PFC ensemble:
@@
$\|\Phi^{(m)}_{\rm Haar}\otimes\mathcal{I}(\rho)-\Phi^{(m)}_{\rm rPFC}\otimes\mathcal{I}(\rho)\|_{\rm tr}\leq O(m/\sqrt{d}),$ (S5)
where $\Phi^{(m)}_{\rm Haar}\otimes\mathcal{I}(\rho)=\mathbb{E}_{U\sim{\rm Haar}}[(U^{\otimes m}\otimes I)\rho(U^{\otimes m}\otimes I)^{\dagger}]$ is the $m$-fold twirling channel for the Haar ensemble (acting only on the $\mathcal{H}_{A}^{\otimes m}$ factor of the Hilbert space), and $\Phi^{(m)}_{\rm rPFC}$ is the same for the random PFC ensemble.
@@
The moment operators for the two ensembles $\mathcal{E}_{\rm rMPS^{\prime}}$, $\mathcal{E}_{\rm rMPS}$ can be written as
@@
$\rho^{(m)}_{\mathcal{E}_{\rm rMPS^{\prime}}}$ $=\Phi^{(m)}_{\rm rPFC}|_{[L,L+\nu]}\circ\Phi^{(m)}_{\rm rPFC}|_{[L-1,L-1+\nu]}\circ\cdots\circ\Phi^{(m)}_{\rm rPFC}|_{[1,\nu+1]}(\rho_{0}),$ (S7)
@@
$\|\rho^{(m)}_{\mathcal{E}_{\rm rMPS}}-\rho^{(m)}_{\mathcal{E}_{\rm rMPS^{\prime}}}\|_{\rm tr}\leq L\epsilon,$ (S8)
with $\epsilon$ the approximation error of the random PFC ensemble on $\nu+1$ qubits: $\epsilon\leq O(m/\sqrt{\chi})$, where $\chi=2^{\nu}$ is the MPS bond dimension. We will prove the statement by induction.
  • ($L=1$): The whole system consists of $\nu+1$ qubits. We have $\|\Phi^{(m)}_{\rm Haar}(\rho_{0})-\Phi^{(m)}_{\rm rPFC}(\rho_{0})\|_{\rm tr}\leq\epsilon$.
@@
    $\rho_{2}$ $=\Phi^{(m)}_{\rm rPFC}|_{[L-1,L-1+\nu]}\circ\cdots\circ\Phi^{(m)}_{\rm rPFC}|_{[1,\nu+1]}(\rho_{0}).$ (S10)
@@
    $\|\Phi^{(m)}_{\rm Haar}|_{[L,L+\nu]}(\rho_{1})-\Phi^{(m)}_{\rm rPFC}|_{[L,L+\nu]}(\rho_{2})\|_{\rm tr}$ $\leq\|\Phi^{(m)}_{\rm Haar}|_{[L,L+\nu]}(\rho_{1})-\Phi^{(m)}_{\rm rPFC}|_{[L,L+\nu]}(\rho_{1})\|_{\rm tr}$
    $\qquad+\|\Phi^{(m)}_{\rm rPFC}|_{[L,L+\nu]}(\rho_{1})-\Phi^{(m)}_{\rm rPFC}|_{[L,L+\nu]}(\rho_{2})\|_{\rm tr}.$ (S11)
    The first term on the rhs is $\leq\epsilon$ by direct application of Eq. (S5). The second term, by monotonicity of the trace distance under quantum channels, is $\|\Phi^{(m)}_{\rm rPFC}|_{[L,L+\nu]}(\rho_{1}-\rho_{2})\|_{\rm tr}\leq\|\rho_{1}-\rho_{2}\|_{\rm tr}$; but (up to tensoring with a trivial state |$\outerproduct{0}{0}^{\otimes m}$) $\rho_{1}-\rho_{2}$ is the difference between the two moment operators at step $L-1$, which by hypothesis is $\leq(L-1)\epsilon$. The result follows.
We remark that this proof did not use the MPS structure specifically, but only the isometric structure of the tensor network (i.e. the ability to write it as a sequential application of unitary gates). As long as the network consists of ${\rm poly}(N)$ gates, the error incurred by replacing Haar-random gates by random PFC gates remains small, $O({\rm poly}(N)m/\sqrt{\chi})$, see Sec. S3.2
@@
Here we modify the proof presented in the main text for MPS to general isometric tensor network architectures. We define the state ensembles $\mathcal{E}_{\rm pTNS}$ (made of pseudorandom PFC gates), $\mathcal{E}_{\rm rTNS^{\prime}}$ (made of random PFC gates) and $\mathcal{E}_{\rm rTNS}$ (made of Haar-random gates) in the same ways as $\mathcal{E}_{\rm pMPS}$, $\mathcal{E}_{\rm rMPS^{\prime}}$ and $\mathcal{E}_{\rm rMPS}$, just replacing the MPS architecture with a general tensor network state (TNS).
@@
    Unitary vertices must have both incoming and outgoing edges, and must respect the condition $\prod_{i\in{\rm incoming}}\chi_{i}=\prod_{i\in{\rm outgoing}}\chi_{i}$ (required by unitarity).
@@
    Input vertices have only outgoing edges. They represent the injection of auxiliary qudits in the system. We restrict for simplicity to product states |$\ket{0}$, which have one outgoing edge, and Bell pairs |$\ket{\psi_{\rm Bell}}=(1/\sqrt{\chi})\sum_{j=0}^{\chi}\ket{jj}$, which have two outgoing edges of the same bond dimension $\chi$. Both types of input vertices are illustrated in Fig. 2(a,c).
@@
    Output vertices have only one incoming edge and no outgoing edges. They represent degrees of freedom of the final tensor network state, or “dangling legs” in standard tensor network notation.
In analogy with the MPS case, we define the number of output vertices as $L$. We also define $N$ as the effective system size (i.e. number of qubits) via $N=\log_{2}(D)$, with $D=\prod_{\ell\in{\rm output}}\chi_{\ell}$ the total Hilbert space dimension. $N$ need not be an integer. We assume that the total number of unitary vertices in the graph is $n_{U}\leq O({\rm poly}(N))$, which ensures the states’ efficient preparability.
@@
We can repeat the same argument as in Sec. S2, but instead of $O(L\epsilon)$ as the upper bound on the trace distance [Eq. (S8)] we obtain $O(\sum_{u\in{\rm gates}}\epsilon_{u})$, since in this case each gate may act on a Hilbert space of different dimension. We have $\epsilon_{u}\leq O(m^{2}/d_{u})$, with $d_{u}=\prod_{\ell\in{\rm outgoing}(u)}\chi_{\ell}$, and thus
$\|\rho^{(m)}_{\mathcal{E}_{\rm rTNS^{\prime}}}-\rho^{(m)}_{\mathcal{E}_{\rm rTNS}}\|_{\rm tr}\leq O(n_{U}m/\sqrt{d_{\rm min}}),\qquad d_{\rm min}=\min_{u}\prod_{\ell\in{\rm outgoing}(u)}\chi_{\ell}.$ (S12)
Here $n_{U}$ is the total number of unitary gates in the network, again assumed to be $O({\rm poly}(N))$. If all the $\chi_{i}$ are super-polynomial in $N$, then $d_{\rm min}=\omega({\rm poly}(N))$ and the two ensembles are indistinguishable in polynomial time.
@@
Here we generalize Lemma 2 to the more general class of tensor network states (TNS) beyond MPSs. We aim to prove that the ensemble of TNSs built out of Haar-random gates is indistinguishable from the Haar-random ensemble on the whole Hilbert space. In analogy with the MPS case, we can compute the $m$-copy distinguishability, perform the Haar averages, and obtain a modified graph corresponding to the partition functions in Fig. 2(b,d). The new graph is undirected, weighted and labeled—each edge has one of two labels, that we denote by ‘solid’ and ‘dashed’ in keeping with Fig. 1 and Fig. 2. The new graph (interaction graph of the partition function) is built from the old one (the tensor network diagram) according to these rules:
  • All edges between two unitary vertices become ‘solid’ and keep their weight $\chi_{i}$.
@@
This graph serves as the interaction graph for an $S_{m}$-valued magnet’s partition function: each vertex hosts a spin $\alpha\in S_{m}$, each ‘solid’ edge connecting two spins $\alpha_{i,j}$ contributes a Boltzmann weight $\chi_{\ell}^{|\alpha_{i}\alpha_{j}^{-1}|}$, each ‘dashed’ edge connecting two spins $\alpha_{i,j}$ contributes $\textsf{Wg}_{d_{u}}(\alpha_{i}\alpha_{j}^{-1})$, with $\chi_{\ell}$ and $d_{u}$ the edge weights as defined above. In addition, there is an overall prefactor of $\prod_{\ell\in{\rm Bell}}\chi_{\ell}^{-m}$, with the product running over Bell pair input states in the original graph, coming from the normalization prefactor in |$\outerproduct{\psi_{\rm Bell}}{\psi_{\rm Bell}}^{\otimes m}$.
@@
where we used the results of Sec. S1 to approximate the Weingarten functions, and restricted the product over edges to edges that did not originate from Bell pairs. Let us now think about this product in terms of the original (tensor network) graph. Writing $d_{u}=\prod_{\ell\in{\rm outgoing}(u)}\chi_{\ell}$ we may view Eq. (S13) as a product over edges of the graph. Edges between input vertices (whether product or Bell-pair) and unitaries do not appear in the product; edges connecting two unitaries cancel (they contibute a $\chi_{\ell}^{m}$ to the first factor and a $\chi_{\ell}^{-m}$ to the second); edges between unitaries and output vertices contribute a $\chi_{\ell}^{-m}$ factor. In all, the result is $(\prod_{\ell\in{\rm output}}\chi_{\ell})^{-m}=D^{-m}$, with $D$ the physical Hilbert space dimension. The error terms stem only from the Weingarten functions and can be bounded above by $|[1+O(m^{2}/d_{u})]^{n_{U}}-1|$ with $n_{U}$ the number of unitaries. Overall this gives